### v1

# QCS:Feature Refining from Quadruplet Cross Similarity for Facial Expression Recognition

Abstract
On facial expression datasets with complex and numerous feature types, where the significance and dominance of labeled features are difficult to predict, facial expression recognition(FER) encounters the challenges of inter-class similarity and intra-class variances, making it difficult to mine effective features. We aim to solely leverage the feature similarity among facial samples to address this. We introduce the Cross Similarity Attention (CSA), an input-output position-sensitive attention mechanism that harnesses feature similarity across different images to compute the corresponding global spatial attention. Based on this, we propose a four-branch circular framework, called Quadruplet Cross Similarity (QCS), to extract discriminative features from the same class and eliminate redundant ones from different classes synchronously to refine cleaner features. The symmetry of the network ensures balanced and stable training and reduces the amount of CSA interaction matrix. Contrastive residual distillation is utilized to transfer the information learned in the cross module back to the base network. The cross-attention module exists during training, and only one base branch is retained during inference. our proposed QCS model outperforms state-of-the-art methods on several popular FER datasets, without requiring additional landmark information or other extra training data. The code is available at https://github.com/birdwcp/QCS.
Introduction
Facial expression is crucial for machines to comprehend human emotions. Facial expression recognition (FER) holds vast research potential and application worth in human-computer interaction, emotion analysis, and intelligent healthcare. Over decades, deep learning advancements in computer vision have propelled significant progress in facial expression recognition.
However, FER is still a complex task facing many difficulties such as Inter-class similarity and intra-class variances. Faces with varying expressions retain the common features that constitute a human face, yet same expressions can vary widely in pose, identity, illumination, etc. Traditional object classification methods struggle with complex face datasets due to assumptions of labeled feature dominance. Dividing datasets by labels for training can result in the learning of unlabeled redundant features if they’re non-randomly distributed, impacting classification. Fig. 1 illustrates an extreme case of this impact. In reality, interference features are varied, numerous, and unpredictable, exerting varying degrees of influence.
Figure 1: Assuming that unlabeled gender features and labeled expression features both exhibit similar distributions in the training set, leading to the non-dominance of significant expression features. It is difficult to explain whether the trained model is an expression classifier or of another type.
How to extract discriminative features and how to eliminate the interference of redundant features are the core issues of FER. Using self-attention, some deep learning methods (Wang et al. 2020; Xue, Wang, and Guo 2021; Xue et al. 2022; Ma, Sun, and Li 2021) focus on key features, enhancing model discrimination. However, model performance hinges on the amount and quality of training data. Some methods (Xie, Hu, and Wu 2019; Zhang et al. 2021; Cai et al. 2021; Ruan et al. 2021, 2022) study separating redundant from expression features using GANs or transformers. They extract expression information via decomposition and reconstruction, but often need redundant feature labels, such as age, gender, race, or rely on identity recognition modules. Additionally, interference decoupling FER (Li, Deng, and Du 2017) relies on one or several interference feature labels and can only address limited and specific interference factors. However, interference features are often unclear in type and quantity, and these methods that rely on labeled data often fail to achieve good performance when the types, quantities, and quality of labeled data are limited. Furthermore, when the expression feature labels are ambiguous (She et al. 2021; Zhang, Wang, and Deng 2021), it is unrealistic to clearly label all interference features.
Unlike the aforementioned supervised methods, we introduce a straightforward yet potent approach. It harnesses similarities among same-class image pairs to extract discriminative label features and employs similarities in cross-class image pairs to separate unlabeled, redundant interference features, to alleviate the problem of relying on annotated features to be dominant in the training set.
Firstly, We propose Cross Similarity Attention (CSA) tailored for mining fine-grained feature similarity between different images. It distinguishes from NLP Transformer’s cross-attention(Vaswani et al. 2017) for language correlations and Vision Transformer’s self-attention(Dosovitskiy et al. 2021) for intra-image correlations. CSA calculates Euclidean distance similarity between fine-grained features of images, focusing solely on the importance of the features in the spatial dimension of the queried image(key). This compensates for ViT’s matrix multiplication-based attention, which lacks input-output positions correlation in cross-image similarity.
By leveraging CSA for interaction module construction, we enhance the distinct shared features between images within the same class. Inspired by triplet loss (Schroff, Kalenichenko, and Philbin 2015), we apply the interaction module to extract and diminish similar features across different class image pairs, thereby separating redundant and interfering features. To ensure equal feature refinement across input branches, we introduce a four-branch configuration, with each class contributing two images. This forms a centrally symmetric closed-loop network utilizing four sets of interaction modules for stable joint training. This approach simultaneously minimizes intra-class feature differences and maximizes inter-class differences, mimicking the effect of triplet loss at the fine-grained feature level. During inference, only one image is input, requiring the removal of the multi-branch interaction module, retaining a single base network. To effectively integrate the refined information from cross modules, we introduce residual connections that minimize gradient conflicts in feature maps, creating a self-distillation-like structure.
In summary, our main contributions are summarized as follows:
  • We introduce Cross Similarity Attention (CSA), a tailored mechanism for mining fine-grained feature similarities across different images, utilizing a spatially dimensioned and input-output position-sensitive attention approach.
  • We propose Quadruplet Cross Similarity (QCS), a four-branch, centrally symmetric closed-loop joint training framework based on CSA. It refines features by mining discriminative features within the same class and separating interfering features across different classes synchronously.
  • Our method achieves state-of-the-art performance on several FER datasets without requiring additional landmark information or pre-training on FER external datasets.
Figure 2: The framework of Quadruplet Cross Similarity Network. Joint training is performed on 4 classifiers ${{Cls}_{base}}$ based on a weight-shared backbone and 4 classifiers ${{Cls}_{cross}}$ enhanced by cross-module, with only the red branch retained during inference. Anchor and pos are in the same category, so are neg and neg2. The matrix ${SD}$ performs attention refinement on matrix ${S}$ and matrix ${D}$ by rows or columns.
RELATED WORK
Facial Expression Recognition
With the development of deep learning, FER has become increasingly popular in recent years, as more and more researchers focus on human-computer interaction. RAN (Wang et al. 2020) proposes a Region Attention Network that adaptively captures the importance of different region features. SCN(Ma, Sun, and Li 2020) introduces the Self-Cure Network, which suppresses uncertainty in facial expression datasets by excluding samples with ambiguous annotations. DMUE(She et al. 2021) leverages image pair similarity comparisons to mine real label confidence and potential distribution information. RUL(Zhang, Wang, and Deng 2021) computes relative uncertainty through image comparisons, addressing uncertainty arising from blurry images and inaccurate labels.
Transfer(Xue, Wang, and Guo 2021) is the first attempt to apply Vision Transformer (ViT) to adaptively characterize the relations between different facial parts. APViT(Xue et al. 2022) proposes two Attentive Pooling modules with Vision Transformer to focus on the most discriminative features and discard the less relevant ones. POSTER V2(Mao et al. 2023), LA-Net(Wu and Cui 2023), and S2D(Chen et al. 2024) enable effective collaboration between facial landmark features and direct image features.
Attention Mechanisms
Self-attention primarily captures dependencies within a sequence to focus on important information. SENet(Hu, Shen, and Sun 2018) introduced a plug-and-play channel attention module. BAM(Park et al. 2018) and CBAM(Woo et al. 2018) introduced spatial attention modules based on convolutions. However, these convolution-based spatial attention mechanisms can only perform local information interaction. The Non-Local Block(Wang et al. 2018) can capture long-range dependencies between any two positions in images. Subsequently, the Transformer(Vaswani et al. 2017) based on Scaled Dot-Product Attention became the state-of-the-art (SOTA) for natural language processing (NLP) tasks. ViT(Dosovitskiy et al. 2021) was the first to implement a pure transformer structure in the field of computer vision.
Using distinct features for computing interaction relationships constitutes cross-attention. In Transformer, K and V sequences stem from the source encoder, while Q sequence originates from the target decoder. In Few-shot classification tasks, cross-attention in CEC and CAN(Lai et al. 2023; Hou et al. 2019) can extract more discriminative features even when the sample size is limited. Fine-grained image classification methods CIN(Gao et al. 2020) and API(Zhuang, Wang, and Qiao 2020) propose channel cross-attention, which is used to extract channel similarities between image pairs. CrossViT(Chen, Fan, and Panda 2021) unifies self-attention features across different scales by fusing information from small-patch and large-patch tokens. In multimodal feature fusion(Wei et al. 2020; Nagrani et al. 2021; Zhang et al. 2023), cross-attention is used to exchange information between different modalities through early, middle, or late fusion.
Contrastive Learning
Contrastive learning is a type of Self-Supervised Learning method(Chen et al. 2020; Chen, Xie, and He 2021). Supervised Contrastive Learning(Khosla et al. 2020) introduces labeled and different data of the same class for contrastive learning. Mutual Contrastive Learning(Chuanguang Yang 2022) collaboratively trains a cohort of models from the perspective of contrastive representation learning. Triplet loss(Schroff, Kalenichenko, and Philbin 2015) is a commonly used contrastive learning loss function in face recognition tasks. FEC(Vemulapalli and Agarwala 2019) learns a compact and continuous expression embedding space by comparing the similarity of three images using triplet loss.
METHOD
Cross Similarity Attention
Intuitively, comparing the similarities and differences between images can make it easier to judge the category of an image. This similarity comes not only from high-level semantic features but also from relatively low-level fine-grained features. Based on the non-local block(Wang et al. 2018), we propose a spatial attention called Cross Similarity Attention (CSA) to focus on the similarity between different images. This is significantly different from the Dot-Product Attention in ViT(Dosovitskiy et al. 2021), and it does not require concatenation on spatial dimension of different images like the non-local block.
(a) Self Attention
(b) Cross Attention
(c) Cross Similarity Attention
(d) Similarity Matrix
Figure 3: Different types of attention on feature vectors h×w×c, where h=w=2, and c denotes channels. Blue and yellow represent features from different images, and each element corresponds to a 1×1×c vector. Double-arrowed lines denote the interaction values between corresponding two elements, and thicker lines indicate higher weights. The 4x4 green box in (d) represents the interaction matrix of (c).
Firstly, In standard ViT qkv self-attention (as shown in Fig. 3(a)), the query ${Q}$, key ${K}$, and value ${V}$ are all feature vectors derived from the same image. The output feature position is the same as the query feature position. The weight vectors can reflect the interactions between the input positional features and the output positional features.
However, if the query ${Q}$ comes from the features of one image p, while the key ${K}$ and values ${V}$ are derived from the features of a different image a, as shown in Fig. 3(b), some changes occur. The output positional features(a0, v0) in image ${a}$ are different from the query positional feature in image ${p}$(p0) (this is different from self-attention). The weight ${w}$ reflects the strength of the association with the values ${V}$, but it cannot reflect the strength of the association with the output vector out(t0) at the position in image ${a}$.
Therefore we consider only quantifying the importance of the positions in the queried image (key ${K}$), regardless of where the query ${Q}$ comes from, and using these weights to weight the features at the output positions (values ${V}$) of that image, as shown in Fig. 3(c) 3(d).
Figure 4: Interaction matrix of 7x7xC (H=7, W=7, channel=C) image features in the spatial dimension (HxW=49). (a) cosine similarity SA (b) dot product SA (c) euclidean distances SA (d) euclidean distances CA
Secondly, the self-attention in the standard ViT tends to capture the interactions rather than the similarity between vectors. The $D_{h}$ in ${W_{s}\left({Q,K}\right)=softmax(}{{{Q}{K}^{{T}}}/\sqrt{{D}_{{h}}}}{)}$ is a unified scaling factor. If we emphasize similarity, when calculating self-attention(SA), the vector at a certain spatial position in the same image should have a higher response with itself, as shown in Fig. 4(a) 4(c). The yellow on the diagonal of the cosine similarity or Euclidean distance similarity matrix is more prominent. However, the dot product(Fig. 4(b)) calculation used in ViT weakens the response intensity on the diagonal of the interaction matrix. Meanwhile, the similarity matrix(Fig. 4(d)) between different images based on cross-attention(CA) lacks a highlighted diagonal, where ${\text{S}_{\text{c}\text{s}}\left(\text{Q,K}\right)}_{\text{ij}}\neq{\text{S}_{\text{c}\text{s}}\left(\text{Q,K}\right)}_{\text{ji}}$.
The original non-local block captures similarity among concatenated frames that is of practical significance, utilizing self-attention, but inflates the interaction matrix. In contrast, our CSA(Fig. 3(c)) efficiently computes similarity between different feature maps.
$D_{cs\_pa}\left(Q_{p},K_{a}\right)=\parallel Q_{p}-K_{a}\parallel\in\mathbb{R}^{l\times l}$ (1)
$\begin{split}S_{cs\_pa}\left(Q_{p},K_{a}\right)=D_{cs\_pa}\left(Q_{p},K_{a}\right)_{\max}\\ -D_{cs\_pa}\left(Q_{p},K_{a}\right)\in\mathbb{R}^{l\times l}\end{split}$ (2)
We calculate the similarity interaction matrix using Euclidean distance as shown in Eq. 1 2, ${D_{cs}\left({Q,K}\right)_{ij}=\sqrt{\sum_{c=1}^{C}(Q_{ic}-K_{jc})^{2}}}$, ${Q,K,V\in\mathbb{R}^{l\times c}}$ where ${l=w\times h}$) and separately assess the importance of the queried positions in both images, thereby emphasizing the weights of common features with high similarity across different spatial locations within each image. Inspired by GCNet(Cao et al. 2019), which observed that attention maps calculated for different query positions are nearly identical, we aggregate the interaction matrix by rows or columns.
$S_{cs\_a}\left(Q_{p},K_{a}\right)=\sum_{i=1}^{l}\left(\operatorname{norm}\left(\left(S_{cs\_pa}\left(Q_{p},K_{a}\right)\right)_{i,:}\right)\right)_{ij}\in\mathbb{R}^{1\times l}$ (3)
$A_{cs\_a}\left(Q_{p},K_{a}\right)=\operatorname{softmax}\left(S_{cs\_a}\left(Q_{p},K_{a}\right)\right)^{T}\cdot V_{a}\in\mathbb{R}^{l\times 1}$ (4)
Eq. 3 4 calculate the attention for image ${a}$. Here, ${\operatorname{norm}\left({S_{cs}\left({Q,K}\right)}_{i,:}\right)}$ normalizes the similarity matrix by row using the ${L2}$ norm to eliminate absolute differences in similarity calculated from query features at different positions, then sums the values for the same queried spatial position across columns. Finally, a mutually exclusive softmax function is employed to directly compute the weight distribution along the spatial dimension, where ${\sum_{j=1}^{l}\left(\operatorname{softmax}\left({S_{cs\_a}\left({Q,K}\right)}\right)^{T}\right)_{j}}=1$. Eq. 5 6 symmetrically calculate the attention for image ${p}$.
$S_{cs\_p}\left(Q_{p},K_{a}\right)=\sum_{j=1}^{l}\left(\operatorname{norm}\left(\left(S_{cs}\left(Q_{p},K_{a}\right)\right)_{:,j}\right)\right)_{ij}\in\mathbb{R}^{l\times 1}$ (5)
$A_{cs\_p}\left(Q_{p},K_{a}\right)=\operatorname{softmax}\left(S_{cs\_p}\left(Q_{p},K_{a}\right)\right)\cdot V_{p}\in\mathbb{R}^{l\times 1}$ (6)
Quadruplet Cross Similarity Network
By leveraging Cross Similarity Attention (Section 3), fine-grained features with high similarity between two images belonging to the same category can be easily extracted, and their weights enhanced. The resulting classifier, ${{Cls}_{cross}}$ , outperforms the backbone-based ${{Cls}_{base}}$ in discrimination. This approach is termed the Dual Cross Similarity (DCS) Network.
Figure 5: cross similarity heatmap.yellow represents higher similarity, and blue represents greater distance, i.e., lower similarity. The cross similarity vectors ${{S}_{cs\_p}(1\times 49)}$ and ${{S}_{cs\_a}(49\times 1)}$ are reshaped to ${7\times 7}$ heatmaps.
However, merely relying on Cross Similarity Attention for feature weighting in same-category image pairs has limitations. Highly similar features may not be discriminative, e.g., flat edge areas in most face images. Unlabeled features like gender, skin tone, etc., may exhibit similarity between images of the same or different categories. Leveraging distribution differences with labeled features via cross similarity attention can mine redundant interfering features, separating them from similar same-category features, refining cleaner features. The cross similarity heatmaps between the anchor image and images of the same or different categories are shown in Fig. 5.
To achieve the joint cross similarity matrice from both same and different categories through addition or multiplication operations, we have adopted Euclidean distance instead of cosine similarity. We directly utilize the distance matrix D of different categories for coarse alignment with the similarity matrix S of the same category in terms of their numerical values, as shown in Eq. 7. By applying L2 normalization to the interaction matrix row-wise or column-wise, as in Eq. 8, we further achieve fine alignment.
$\begin{split}D_{cs\_na}\left({Q_{n},K_{a}}\right)=D_{cs\_na}\left({Q_{n},K_{a}}\right)\\ -D_{cs\_na}\left({Q_{n},K_{a}}\right)_{min}\in\mathbb{R}^{l\times l}\end{split}$ (7)
$D_{cs\_a}\left({Q_{n},K_{a}}\right)={\sum_{i=1}^{l}\left(norm\left({D_{cs\_na}\left({Q_{n},K_{a}}\right)}_{i,:}\right)\right)_{ij}}~{}\in\mathbb{R}^{1\times l}$ (8)
We found that adding matrices(Eq. 9) led to better performance than multiplying them. ${\gamma}$ is a hyperparameter and ${\theta}$ is a learnable parameter. The nonlinear activation function ${\tanh(\gamma\cdot\theta)}$ is used to limit the value to (-1,1), and more effective gradient propagation is obtained near the initial value ${\theta=0}$, enabling adaptive enhancement or suppression of matrix D on S. Finally, the vector ${{SD}_{cs\_a}}$ is used to replace ${{S}_{cs\_a}}$ in Eq. 4.
$\begin{split}{SD}_{cs\_a}\left({Q_{p},K_{a},Q_{n}}\right)=S_{cs\_a}\left({Q_{p},K_{a}}\right)\\ +\left({1+{\tanh(\gamma\cdot\theta)}}\right)\cdot D_{cs\_a}\left({Q_{n},K_{a}}\right)\in\mathbb{R}^{1\times l}\end{split}$ (9)
To enhance similar features within the same category while emphasizing features with large distances between different categories. This is similar to the triplet loss(Schroff, Kalenichenko, and Philbin 2015), which reduces the feature differences within the same category and simultaneously increases the feature differences between different categories. However, if only the anchor-positive-negative triplet is introduced, the negative cannot construct a symmetric cross network with the anchor or positive due to the absence of a paired image from the same category. We have attempted to freeze the negative branch while only utilizing its redundant interfering features separated from the anchor branch, but there is a significant performance gap between this network and the joint training network, possibly due to the issue of gradient stability.
Therefore, we added another negative image, ingeniously forming the four branches into a centrally symmetric closed-loop network, as shown in Fig. 2. This approach allows us to calculate only two S matrices for the same category and two D matrices for different categories, with each interaction matrix being utilized by two adjacent images simultaneously.The features obtained from the backbone without the class token go through an FC layer to produce vectors ${QK}$ and ${V}$. When calculating the interaction matrices with the features of two adjacent images, the weights of ${QK}$ are shared and act as either ${Q}$ query or ${K}$ key. Consequently, the interaction matrices exhibit symmetry, where ${(Q2K1)^{T}=Q1K2}$ and ${(Q3K4)^{T}=Q4K3}$. When ${SD}$ fuses the matrices ${S}$ and ${D}$ by rows or columns, it reduces the need for two interactive calculations. The Quadruplet Cross Similarity(QCS) Network offers computational efficiency while achieving stability in multi-branch joint training.
Contrastive Residual Distillation
The training loss curve in Fig. 6 shows a significant performance gap between ${{Cls}_{base}}$ and the enhanced ${{Cls}_{cross}}$ by the cross module. During inference, only one ${{Cls}_{base}}$(red branch in Fig. 2) is retained to predict results, posing a challenge in effectively transferring cross module information into the inference network.
A straightforward approach is to introduce residual connections between ${{Cls}_{base}}$ and ${{Cls}_{cross}}$ for direct gradient feedback. We explored diverse connection methods, such as element-wise addition followed by Global Average Pooling (GAP) and Bilinear pooling (BP) to fuse and pool the feature vectors. Additionally, we considered treating ${{Cls}_{cross}}$ as the teacher model and ${{Cls}_{base}}$ as the student model, introducing a Kullback-Leibler (KL) divergence loss or L2 loss to construct a self-distillation learning(Zhang et al. 2019; Romero et al. 2015; Liu, Tang, and Wu 2020). We further integrated a ViT module into ${{Cls}_{cross}}$ to align with ${{Cls}_{base}}$, minimizing gradient conflicts on feature maps. We term the transfer of information from interactive contrastive learning in the cross module to the base network as Contrastive Residual Distillation.
$\mathcal{L}_{total}={\sum_{i=1}^{C}\mathcal{L}_{{ce}_{base\_i}}}+\lambda_{1}\cdot\mathcal{L}_{{ce}_{cross\_i}}+\left[(\lambda_{2}\cdot\mathcal{L}_{{kl}_{i}})or(\lambda_{3}\cdot\mathcal{L}_{{2}_{i}})\right]$ (10)
The joint training loss, utilizing cross entropy for all classifiers, is defined in Eq. 10. C set to 2 for DCS and 4 for QCS, The [] indicates an optional element.
EXPERIMENT
Datasets
RAF-DB(Li, Deng, and Du 2017) is a real-world affective face database which contains 29,672 facial images labeled with basic or compound facial expressions. In the experiment, 15,331 images with 7 basic expressions(i.e. surprise, fear, disgust, happiness, sadness, anger, neutral) are choosed , of which 12271 are used for training and 3068 for testing.
FERPlus(Fer 2016) is extended from FER2013(consists of 28,709 training images and 3,589 test images) and relabeled with 8 expression labels (neutral, happiness, surprise, sadness, anger, disgust, fear, contempt) by 10 taggers. In the experiment, we follow the code(Fer 2016) to handle majority voting, which keeps a single label for each image and ignores the samples with unknown or NF labels, then measure prediction accuracy on the test set.
AffectNet(Mollahosseini et al. 2017) is a large-scale FER dataset in the wild, which contains about 440,000 manually annotated facial images. We utilized AffectNet-7/8 with 7/8 expressions, comprising 283,901/287,651 training and 3,500/4,000 validation images. Balance-sampling(Imb 2019) is used since the training set exhibits a long-tailed distribution, while the validation set is balanced.
Implementation Details
We adopted POSTER++ removing the landmark branch as the baseline and constructed our model upon it. All face images are resized to 224 × 224 pixels for training and testing. We use IR50(Deng et al. 2019), which is pretrained on the Ms-Celeb-1M dataset(Guo et al. 2016), as the backbone. We employ an Adam optimizer with Sharpness Awareness Minimization (SAM)(Foret et al. 2020) to train the model(200 epochs for RAF-DB, 100 epochs for FERPlus and 40 epochs for AffectNet). For the DCS, the batch size is set to 48, the learning rate is initialized as 9e-6 and an exponential decay learning rate scheduling with a gamma of 0.98 is employed. For the QCS, the batch size and the initial learning rate are reduced to around half. It is trained end-to-end on a single Nvidia RTX3090 GPU via Pytorch.
(a) row 4
(b) row 5
(c) row 6
(d) row 7
(e) D
(f) MSE
Figure 6: The training loss curves of ${Cls_{base}}$ and ${Cls_{cross}}$. (a)(b)(c)(d) correspond to the models in rows 4-7 of Table 1, respectively. (e)(f) correspond to the QCS, displaying only the anchor and pos branches.
Ablation Analysis
Initially, we trained a classification network utilizing the IR50+ViT backbone as a baseline. Then, a DCS Network with shared weight branches was constructed on the backbone to evaluate the performance impact of each component in our method through ablation studies on RAF-DB and FERPlus. Each experiment is repeated five times, then the maximum accuracy is reported. Unless otherwise stated, the experimental settings adopt the default configurations outlined above.
[…]
ViT GAP ViT
1 92.05 91.06
2 91.98 90.83
3 91.95 90.93
4 92.14 91.02
5 92.24 91.25
6 92.57 91.22
7 92.21 90.93
[…]
Evaluation of cross attetions.
Row 1 of Table 1 presents the accuracy of the baseline, while rows 3 and 4 show that our CSA outperforms SDPA in the comparison. For a fair comparison, all structures remain consistent except for varying core interaction approaches, with Multi-Head Attention and class token removed for maximum structural simplicity. Features (7×7×c) extracted from 3 abstraction levels of the backbone interact individually and are concatenated on the spatial dimension. CSA avoids computing similarities between features of different levels, and SDPA adopts the same approach. The FC and MLP (Multi-Layer Perceptron) of the interaction module share weights across the interaction branches.
Evaluation of residual connection.
We observe the performance gap between ${Cls_{base}}$ and ${Cls_{cross}}$ through their training loss curves(Fig. 6), yet we are unable to obtain the accuracy of ${Cls_{cross}}$ during the inference stage. The model accuracy in row 2 of Table 1 is lower than the baseline(row 1), indicating the gradient differences after multiple layers may have distorted the baseline model’s original feature representation. The improvement from rows 2 to 4 of Table 1 demonstrates the information transfer through residual connections. Rows 4-6 suggest superior performance from minimizing gradient conflicts in connections, that BP retains fine-grained features while ViT module align ${Cls_{cross}}$ with ${Cls_{base}}$. The experiments below adopt the ViT connections.
Evaluation of cross similarity module overfitting.
In Eq. 10, one of $\lambda_{1}$ and $\lambda_{2}$ is set to 1, while the other is set to 0. As shown in Table 1(row 7) and Fig. 6(d), adding the KL divergence loss narrows the losses gap between ${Cls_{base}}$ and ${Cls_{cross}}$, but the accuracy of the model decreases. This phenomenon becomes even more evident in the more complex QCS network using MSE loss, as shown in Fig. 6(f). It is possible that the cross module exhibits overfitting, influencing the base model’s generalization capacity. We set the dropout ratio ranging from 0.2 to 0.5 in the ViT modules to mitigate overfitting during training on AffectNet.
Evaluation of intra-class similarity and inter-class dissimilarity.
[…]
S D SD
RAF-DB 92.18 92.11 92.47
FERPlus 91.12 90.96 91.21
Table 2: Accuracy (%) comparison of different cross matrices.
Table 2 shows S contributes more to improving the model performance compared to D. As can be seen from Fig. 6(e), ${Cls_{cross}}$, which is enhanced by D, does not show loss reduction compared to ${Cls_{base}}$. Fig. 5 shows that the inter-class similarity matrix has more dark blue areas (i.e., more bright yellow highlights in the inter-class distance matrix), indicating more dispersed attention. We infer that although weakening the features with small distances (high similarity) across categories helps separate redundant features, the simultaneously enhanced features with large distances do not necessarily contribute to classification. The SD matrix, resulting from the adaptive fusion of S and D, harnesses the strengths of both S and D to extract more clearly features.
Comparison with the State of the Art
Table 3 compares our best results with the current state-of-the-art methods on RAF-DB, AffectNet-7, and FERPlus. APViT and the other methods listed below are Transformer-based methods, while POSTER++, LA-Net, and S2D incorporate additional landmark information. Our proposed method achieves these results without utilizing extra landmark information or other extra training data, ensuring no inference time increase by removing all interactive modules and retaining a single branch. While QCS effectively separates redundant features, the heightened complexity of the interactive module may pose challenges during training, particularly for AffectNet-7 where over-sampling of underrepresented classes can exacerbate overfitting, ultimately leading to a performance trade-off. Nevertheless, our method obtains competitive results on AffectNet-7 and FERPlus when compared to Landmark-Aware methods, outperforming recent state-of-the-art methods on RAF-DB.
Methods RAF-DB FERPlus AffectNet-7
SCN(2020) 87.03 88.01 60.23
RUL(2021) 88.98 - -
DMUE(2021) 89.42 89.51 63.11
TransFER(2021) 90.91 90.83 65.20
Face2Exp(2022) 88.54 - 64.23
EAC(2022) 89.99 89.64 65.32
APViT(2022) 91.98 90.86 66.91
TAN(2023) 90.87 91.00 66.45
POSTER(2023) 92.05 91.62 67.31
POSTER++(2023) 92.21 - 67.49
S2D(2024) 92.21 91.01 66.42
DCS(Ours) 92.57 91.25 67.29
QCS(Ours) 92.47 91.21 66.91
QCS(Ours)‡ 92.83 91.50 -
Table 3: Performance comparison (%) on RAF-DB, AffectNet-7, and FERPlus. ‡means pre-trained on the AffectNet-8(Acc. 63.9).
Visualization
To visually demonstrate the effectiveness of our method in extracting distinctive features and separating redundant features, we calculated the cross similarity between every pair of the 7 expressions. We utilized the QCS described in Section 3 to simultaneously compute the similarity matrices between the anchor and positive as well as between the anchor and negative. Fig. 7 only displays the similarity heatmap used to calculate the anchor attention, and the heatmaps for positive and negative attention are not shown.
The heatmaps in diagonal positions show that the distinctive features extracted from images of the same category mainly concentrate on the areas of eyes and mouth corners. For example, Happy is mainly reflected in the areas of mouth corners and cheeks, while Disgust mainly in the areas of mouth and nasolabial folds. The exaggerated facial expressions of Anger are reflected in the entire facial features area, and the similar enlargement of eyes and mouths in Surprise and Fear results in similar heatmaps. The heatmaps in non-diagonal positions show the redundant features separated from comparing the images in a certain column (Key) with various expressions. As can be seen from the last row of heatmaps, Neutral, which does not have the facial features of expressions, mainly separates out feature regions from various expressions in the corners of the images.
Figure 7: The confusion matrix of Cross Similarity heatmaps on RAF-DB. The anchor image located above a certain column is used as the key to calculate cross similarity on that anchor. The yellow highlighted area represents high similarity.
CONCLUSION
This paper proposes a feature refining network named Quadruplet Cross Similarity(QCS) for FER. The QCS comprises a closed-loop symmetric network formed by four weight-sharing branches and four cross modules for stable joint training. It extracts discriminative features within the same class and synchronously separate interfering features across different classes. The cross module is based on Cross Similarity Attention (CSA), a novel global spatial attention designed for different image pairs, which is input-output position-sensitive. Extensive experiments on several public datasets indicate that our QCS effectively refines the features into a purer form and achieves state-of-the-art performance. Furthermore, effectively transferring the cross-module’s performance back to the basic inference network while avoiding overfitting and ensuring the stability of the model’s performance remains a crucial task.
Acknowledgments
This work was supported by funding from the National Natural Science Foundation of China (U21B2035 and 82372218).